

********************************************************************
UARS WEEKLY STATUS REPORT - 23 JUL 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE has been performing day only operations since Thursday of last
week in support of the yaw maneuver. We returned to normal operations
on Wednesday. Otherwise, the instrument continues to operate in good
health, taking solar data on all available orbital tracking periods. An
average of 60% of data were returned this week accounting for 353
minutes of solar observing time.
***********
SUSIM
***********
SUSIM was powered off for July 16-21. Otherwise, the experiment
successfully gathered its daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for periods of UARS real-time telemetry or onboard tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2003365 has been reprocessed and archived at the GSFC DAAC. Processing
of HEPSA data is temporarily suspended waiting for additional UARS OA data
processing.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004365. Processing of MEPS data is
temporarily suspended waiting for additional UARS OA data processing.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 July 2004 (Orbit 70202, GMT Day 192) through 16
July 2004 (Orbit 70307, GMT Day 198).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 194/2300 70247 AXIS 1 & 2 OFF
HALOE 195/2335 70262 POWER OFF
PEM 196/2255 70277 POWER OFF
SUSIM 197/2144 70291 POWER OFF
SOLSTICE 197/2252 70292 FULLTIME TO DAY-ONLY
The beta angle ranged between the angles of 32.6 degrees to 04.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 18 July 2004 (DOY 200).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 0.64 degrees C. Additional
battery performance data for each battery on 09 July and 16 July 2004
are provided in Appendix D.
The clock error ranged from 10.1 to -10.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
PEM = 02
HALOE = 01
SUSIM = 01
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
Five (5) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT began testing the new version of the Eclipse ground system
software. There are several DR fixes incorporated in the build, so
testing will take some time. The goal for the testing to be complete
is the end of the July. If all testing is successful the new build
will be released to operations.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 July 2004 (DOY 199). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 July 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
70247 194/2300 04-203 PEM AXIS AXIS 1 & 2 OFF
70252 195/0731 04-204 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
70262 195/2335 04-205 HALOE POWER OFF
70271 196/1418 04-206 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
70277 196/2255 04-207 PEM POWER OFF
70291 197/2144 04-208 SUSIM POWER OFF
70292 197/2252 04-209 SOLSTICE FULLTIME TO DAY-ONLY
70302 198/1452 04-210 MPS YAW: TO VT-LV6
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 16 July 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 16 July 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
70252 195/0736 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
70271 196/1418 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 16 July 2004
09 July 2004 - GMT Day 191 Beta = 36.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.62 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-134.4 +5.15/+3.87 +26.4/-14.4 25.1 1.04 19.7
3 +22.4/-067.2 +3.87/+2.60 +25.6/-14.0 25.1 1.02 19.3
16 July 2004 - GMT Day 198 Beta = 04.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 -05.6/-22.4 +4.51/+3.87 +16.8/-12.0 26.7 1.11 18.8
3 +16.8/-22.4 +3.55/+2.92 +16.0/-11.6 26.7 1.11 18.4
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 16 July 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-043 FOT 70203 192/0155 30824 Data Dropout (18 sec)
04-044 SSPP 70234 194/0229 N/A SSPP File Xfer Error on CMS
04-045 FOT 70258 195/1750 36367 Data Dropout @ White Sands
04-046 FOT 70290 197/2031 36385 Data Dropout Bad Weather @ WS
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-040 SOLSTICE 70140 187/2029 N/A SOLSTICE Microload Problem
04-041 FOT 70167 189/1621 36316 Late Acq. due to Guam equip
04-042 FOT 70177 190/0739 36321 Late Acq. Software prob @ WSGT
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-039 FOT See List July 04 30832 Generic Late Acquisition
04-043 FOT 70203 192/0155 30824 Data Dropout (18 sec)
04-044 SSPP 70234 194/0229 N/A SSPP File Xfer Error on CMS
04-045 FOT 70258 195/1750 36367 Data Dropout @ White Sands
04-046 FOT 70290 197/2031 36385 Data Dropout Bad Weather @ WS